OpenStudy (math_bender_15):

Ahmad's gas tank is 1/5 full. After he buys 14 gallons of gas, it is 7/10 full. How many gallons can Ahmad's tank hold?

OpenStudy (pfenn1):

Let x = number of gallons the tank can hold.\[\frac 15 x+14=\frac{7}{10}x\] Solve for x.